Wildhorse Drilling Company has leased property on which oil has been discovered. The oil wells on this property produced 16,300 barrels of oil during the past year that sold at an average sales price of $61 per barrel. Total oil resources of this property

Wildhorse Drilling Company has leased property on which oil has been discovered. The oil wells on this property produced 16,300 barrels of oil during the past year that sold at an average sales price of $61 per barrel. Total oil resources of this property are estimated to be 272,000 barrels.

The lease provided for an outright payment of $557,600 to the lessor (owner) before drilling could be commenced and an annual rental of $27,710. A premium of 5% of the sales price of every barrel of oil removed is to be paid annually to the lessor. In addition, Wildhorse (lessee) is to clean up all the waste and debris from drilling and to bear the costs of reconditioning the land for farming when the wells are abandoned. The estimated fair value, at the time of the lease, of this clean-up and reconditioning is $35,360.

From the provisions of the lease agreement, compute the cost per barrel for the past year, exclusive of operating costs, to Wildhorse Drilling Company. (Round answer to 2 decimal places, e.g. 4.89.)
Answer #1

Initial Payment: 557,600 / 272,000 = 2.05

Rental: 27,710 / 16,300 = 1.70

Premium: 5% * $61 = 3.05

Reconditioning of Land: 35,360 / 272,000 = 0.13


Total Cost Per Barrel: 6.93

answered by: Orange